4pm 29.pd 1. Evaluate the definite integrals, * dx 2. Use an integral to find the shaded area below etx=h
Answer #1

셔 O I = Stadt = A = - 4 = Ca+a3-27h= \+- 터 (대르 0의 로 2 0여 야일게 =I= (2+4-용)-(-2+)= 6월 4월 는